OpenStudy (anonymous):

Solve the equation by completing the square. Round to the nearest hundredth if necessary. x^2-x=25 (A)4.97, 5.02 (B)5.52, –4.52 (C)25.75, –24.75 (D)5.45, –4.45

OpenStudy (whpalmer4):

Put your equation in the form \[ax^2 + bx =c\]Divide through by \(a\) if \(a \ne 1\) \[x^2-x = 25\]Now take half the coefficient of the \(x\) term, square it, and add to both sides. You can rewrite the left hand side as a perfect square \((x-\frac{1}{2})^2\) at this point.. Take the square root of both sides and solve for the two values of \(x\) (remember the square root has a positive and a negative value!)

OpenStudy (dan815):